OpenStudy (anonymous):

Need Help with One Multiple Choice Question. Part A and B.-- Question Inside

OpenStudy (anonymous):

for the first one its 16 i believe

OpenStudy (yb1996):

h is \[8\sqrt{2}\]

OpenStudy (anonymous):

I attached the wrong one. Here is the correct attachment.

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

part b is a

OpenStudy (yb1996):

A) \[\sqrt{\frac{ 3 }{ 2 }}\]

OpenStudy (yb1996):

whoops, the root isn't over the whole thing, just the three

OpenStudy (anonymous):

Okay thank you I'll fan both of you guys. Can you help with a few more?

OpenStudy (yb1996):

sure

OpenStudy (yb1996):

for part A, sin(35) = x/20 so x = 20sin(35)

OpenStudy (yb1996):

for part b cos(41) = x/11 so, x = 11cos(41)

OpenStudy (yb1996):

for part c arcsin(5/14) is about 21 degrees

OpenStudy (anonymous):

C for the second one?

OpenStudy (yb1996):

Yes

OpenStudy (yb1996):

And A for the second one

OpenStudy (yb1996):

C is for the first one, not the second one

OpenStudy (anonymous):

Here are the last one! I appreciate your help!

OpenStudy (yb1996):

part A arccos(3/5.8) = 59 part B x = 30 part C arccos(0.2546) = 75.25

OpenStudy (anonymous):

omg thank you! I appreciate it

OpenStudy (yb1996):

your welcome :)